OpenStudy (anonymous):

A car starting from rest moves with constant acceleration of 2.0 m/s2 for 10 s, then travels with constant speed for another 10 s, and then finally slows to a stop with constant acceleration of -2.0 m/s2. How far does it travel? PLease answer fast!

OpenStudy (***[isuru]***):

Hi, U can solve it by using 2 methods 1. With the help of a (v-t) graph - [velocity -time] 2. with equations of linear motion Which way do u prefer.... to solve this ?

OpenStudy (anonymous):

x1=1/2 a t^2 x2=(vo)t + x1 x3=-(1/2)at^2 + x2

OpenStudy (anonymous):

I prefer to solve it the fast way. I have 4 different answers. 200m, 300m, 400m, or 500m

OpenStudy (***[isuru]***):

It depends on wt u r more familiar with .... R u familiar with equations than graphs or the other way round?

OpenStudy (anonymous):

graphs please. which one is the answer? is it 200m? 300m? 400m? or 500m?

OpenStudy (***[isuru]***):